# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v2

@@ II-A Antenna Boresight Adjustment
where $\theta_{\mathrm{max}}\in[0,\frac{\pi}{2}]$ is the maximum zenith angle that each RA is allowed to adjust. [^1] To characterize the fundamental performance of the RA-enabled wireless systems, we assume that the zenith and azimuth angles of each RA can be continuously tuned in $[0,\theta_{\mathrm{max}}]$ and $[0,2\pi)$, respectively, while in practice they are usually selected from a finite number of discrete values within $[0,\theta_{\mathrm{max}}]$ and $[0,2\pi)$ for the ease of hardware implementation, respectively. The design of RA-enabled systems with discrete deflection angles will be left for future work.
@@ II-B Channel Model
The effective antenna gain for each RA depends on both the signal arrival/departure angle and antenna directional gain pattern, which characterizes the antenna radiation power distribution in different directions. In this paper, we consider a generic directional gain pattern for each RA as follows [25].
@@
where $(\epsilon,\varphi)$ is a pair of incident angles corresponding to any spatial point with respect to the RA’s current boresight direction as shown in Fig. 2(b), $p$ is the directivity factor that characterizes the beamwidth of the antenna’s main lobe, and $G_{0}=2(2p+1)$ is the maximum gain in the boresight direction (i.e., $\epsilon=0$) to satisfy the law of power conservation.
We consider a scattering environment with $Q$ distributed scatterers, where the position of scatterer $q$ is represented by $\mathbf{u}_{q}\in\mathbb{R}^{3\times 1}$. Based on the Friis Transmission Equation and the directional gain pattern adopted in (5), and assuming the gain of the transmit antenna at each user is normalized as $G_{t}=1$, the channel power gain between user $k$ and RA $n$ can be modeled as [26]
$g_{k,n}(\bm{\theta}_{n})$ $\approx\int_{\mathbb{S}_{n}}\frac{1}{4\pi\|\mathbf{q}_{k}-\mathbf{s}\|^{2}}G_{0}\left(\frac{\mathbf{f}^{T}(\bm{\theta}_{n})(\mathbf{q}_{k}-\mathbf{s})}{\|\mathbf{q}_{k}-\mathbf{s}\|}\right)^{2p}\mathrm{d}\mathbf{s}$
$=\frac{S}{4\pi r_{k,n}^{2}}G_{0}\operatorname{cos}^{2p}(\epsilon_{k,n}),$ (6)
@@
where the integral space $\mathbb{S}_{n}=\left[n_{x}d-\frac{\sqrt{S}}{2},n_{x}d+\frac{\sqrt{S}}{2}\right]\times\left[n_{y}d-\frac{\sqrt{S}}{2},n_{y}d+\frac{\sqrt{S}}{2}\right]$ corresponds to the surface region of RA $n$, and $\operatorname{cos}(\epsilon_{k,n})\triangleq\mathbf{f}^{T}(\bm{\theta}_{n})\vec{\mathbf{q}}_{k,n}$ is the projection between user $k$’s direction vector $\vec{\mathbf{q}}_{k,n}\triangleq\frac{\mathbf{q}_{k}-\mathbf{w}_{n}}{\|\mathbf{q}_{k}-\mathbf{w}_{n}\|}$ and the pointing vector of RA $n$. Similarly, the channel power gain between scatterer $q$ and RA $n$ is modeled as
$m_{q,n}(\bm{\theta}_{n})=\frac{S}{4\pi d_{q,n}^{2}}G_{0}\operatorname{cos}^{2p}(\tilde{\epsilon}_{q,n}),$ (7)
@@
where $d_{q,n}=\|\mathbf{u}_{q}-\mathbf{w}_{n}\|$ is the distance between scatterer $q$ and RA $n$, and $\operatorname{cos}(\tilde{\epsilon}_{q,n})\triangleq\mathbf{f}^{T}(\bm{\theta}_{n})\vec{\mathbf{u}}_{q,n}$ is the projection between scatterer $q$’s direction vector $\vec{\mathbf{u}}_{q,n}\triangleq\frac{\mathbf{u}_{q}-\mathbf{w}_{n}}{\|\mathbf{u}_{q}-\mathbf{w}_{n}\|}$ and the pointing vector of RA $n$.
@@
For the multi-path channel between RA $n$ and user $k$, by considering the geometric near-field propagation, the line-of-sight (LoS) channel component $h_{k,n}^{\mathrm{LoS}}(\bm{\theta}_{n})$ and the non-LoS (NLoS) channel component $h_{k,n}^{\mathrm{NLoS}}(\bm{\theta}_{n})$ can be separately modeled by [27, 28]
$h_{k,n}^{\mathrm{LoS}}(\bm{\theta}_{n})$ $=\sqrt{g_{k,n}(\bm{\theta}_{n})}e^{-j\frac{2\pi}{\lambda}r_{k,n}},$ (8)
$h_{k,n}^{\mathrm{NLoS}}(\bm{\theta}_{n})$ $=\sum_{q=1}^{Q}{\frac{\sqrt{\sigma_{q}m_{q,n}(\bm{\theta}_{n})}}{t_{k,q}}}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+t_{k,q})+j\chi_{q}},$ (9)
@@
where $\sigma_{q}$ represents the radar cross section (RCS) of scatterer $q$, modeled as an independent and identically distributed (i.i.d.) positive random variable, $\chi_{q}$ represents the phase shift introduced by scatterer $q$, and $t_{k,q}=\|\mathbf{q}_{k}-\mathbf{u}_{q}\|$ denotes the distance between user $k$ and scatterer $q$. Thus, by superimposing the LoS and NLoS channel components, the overall multi-path channel between user $k$ and the BS is given by
$\mathbf{h}_{k}(\bm{\Theta})=\mathbf{h}_{k}^{\mathrm{LoS}}(\bm{\Theta})+\mathbf{h}_{k}^{\mathrm{NLoS}}(\bm{\Theta}),$ (10)
where $\bm{\Theta}\triangleq[\bm{\theta}_{1},\bm{\theta}_{2},\dots,\bm{\theta}_{N}]\in\mathbb{R}^{2\times N}$ is the deflection angle matrix of all RAs, $\mathbf{h}_{k}^{\mathrm{LoS}}(\bm{\Theta})\triangleq[h_{k,1}^{\mathrm{LoS}}(\bm{\theta}_{1}),h_{k,2}^{\mathrm{LoS}}(\bm{\theta}_{2}),\dots,h_{k,N}^{\mathrm{LoS}}(\bm{\theta}_{N})]^{T}$ and $\mathbf{h}_{k}^{\mathrm{NLoS}}(\bm{\Theta})\triangleq[h_{k,1}^{\mathrm{NLoS}}(\bm{\theta}_{1}),h_{k,2}^{\mathrm{NLoS}}(\bm{\theta}_{2}),\dots,h_{k,N}^{\mathrm{NLoS}}(\bm{\theta}_{N})]^{T}$ are the LoS and NLoS channel components between user $k$ and the BS, respectively.
@@ II-C Min-SINR Maximization Problem
$\mathbf{y}=\sum_{k=1}^{K}{\mathbf{h}_{k}(\bm{\Theta})\sqrt{P_{k}}s_{k}}+\mathbf{n},$ (11)
where $P_{k}$ and $s_{k}$ are the transmit power and information-bearing signal of user $k$, respectively, and $\mathbf{n}$ is the additive white Gaussian noise (AWGN) vector, following the zero-mean CSCG distribution with power $\sigma^{2}$, i.e., $\mathbf{n}\sim\mathcal{N}_{c}(\mathbf{0},\sigma^{2}\mathbf{I}_{N})$. Upon receiving $\mathbf{y}$, the BS applies a linear receive beamforming vector $\mathbf{v}_{k}^{H}\in\mathbb{C}^{1\times N}$ with $\|\mathbf{v}_{k}\|=1$ to extract the signal of user $k$, i.e.,
$y_{k}=\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\bm{\Theta})\sqrt{P_{k}}s_{k}+\sum_{j\neq k}{\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\bm{\Theta})\sqrt{P_{j}}s_{j}}+\mathbf{v}_{k}^{H}\mathbf{n}.$ (12)
Accordingly, the SINR for decoding the information from user $k$ is given by
$\gamma_{k}=\frac{\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\bm{\Theta})|^{2}}{\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\bm{\Theta})|^{2}}+1},$ (13)
In this paper, we aim to maximize the minimum SINR among all the users by jointly optimizing the receive beamforming matrix $\mathbf{V}\triangleq[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{K}]$ and deflection angle matrix $\bm{\Theta}$ of all RAs, subject to their zenith angle constraints in (4). Thus, the optimization problem is formulated as [^2] To evaluate the theoretical performance gain brought by the RA, we assume in this paper that the channel state information (CSI) of all channels involved is perfectly known at the BS. In practice, conventional channel estimation relies on correlating the received signal with a known pilot sequence transmitted from the users (e.g., least-squares (LS) estimation) can be applied at the BS to obtain the required CSI.
@@
s.t. $0\leq\theta_{\mathrm{z},n}\leq\theta_{\mathrm{max}},\;\forall n,$ (14b)
$\|\mathbf{v}_{k}\|=1,\;\forall k.$ (14c)
@@ III Single-User Case with Free-Space Propagation
In this section, we consider the single-user and free-space propagation setup, i.e., $K=1$ and $Q=0$, to draw essential insights into (P1). Thus, the channel modeled in (10) reduces to $\mathbf{h}_{1}(\bm{\Theta})=\mathbf{h}_{1}^{\mathrm{LoS}}(\bm{\Theta})$. In this case, no inter-user interference is present, and thus (P1) is simplified to (by dropping the user index)
@@ III-A Optimal Closed-Form Solution
For any given RA deflection angle matrix $\bm{\Theta}$ in the single-user case, it is known that the MRC beamformer is the optimal receive beamforming solution to problem (P2) [29], i.e., $\mathbf{v}_{\mathrm{MRC}}=\frac{\mathbf{h}^{\mathrm{LoS}}(\bm{\Theta})}{\|\mathbf{h}^{\mathrm{LoS}}(\bm{\Theta})\|}$. Thus, substituting $\mathbf{v}_{\mathrm{MRC}}$ into (15a) yields the following SNR expression,
$\gamma=\bar{P}\|\mathbf{h}^{\mathrm{LoS}}(\bm{\Theta})\|^{2}=\frac{\bar{P}S}{4\pi}\sum_{n=1}^{N}{\frac{G_{0}\operatorname{cos}^{2p}(\epsilon_{n})}{r_{n}^{2}}}.$ (16)
Exploiting the structure in (16), problem (P2) can be decomposed into $N$ subproblems, each of which independently optimizes the deflection angle vector of one RA. For RA $n$, the corresponding subproblem is given by
@@
s.t. $0\leq\theta_{\mathrm{z},n}\leq\theta_{\mathrm{max}},$ (17b)
where the constant term is omitted in (17a). Since $p\geq 0$, by maximizing the projection between the unit vector $\mathbf{f}(\bm{\theta}_{n})$ and $\vec{\mathbf{q}}_{n}$, the optimal solution to problem (P3) is obtained as
$\theta_{\mathrm{z},n}^{\star}$ $=\mathrm{min}\left\{\operatorname{arccos}\left(\vec{\mathbf{q}}_{n}^{T}\mathbf{e}_{3}\right),\theta_{\mathrm{max}}\right\},$ (18a)
@@
where $\mathbf{e}_{1}=[1,0,0]^{T}$, $\mathbf{e}_{2}=[0,1,0]^{T}$, and $\mathbf{e}_{3}=[0,0,1]^{T}$.
Figure 3: Comparison of array directional gain patterns for RA and fixed-antenna arrays with $p=$ 4 and $N_{x}=N_{y}=$ 25.
According to the optimal RA deflection angles in (18), it can be inferred that each RA prefers to adjust its deflection angles to tune its antenna boresight towards the user. This is expected since the BS can achieve the maximum array gain $NG_{0}$ when the boresight direction of each RA is aligned with the user direction, i.e., $\mathbf{f}(\bm{\theta}_{n})=\frac{\mathbf{q}-\mathbf{w}_{n}}{\|\mathbf{q}-\mathbf{w}_{n}\|}$. To demonstrate the essential changes introduced by the proposed RA architecture, based on the directional gain pattern of each antenna element given in (5), we compare the array directional gain patterns of the RA and fixed-antenna arrays in Fig. 3. For the RA array, the boresights of all RAs are adjusted to maximize the SNR in the user direction with $\psi=\frac{\pi}{2}$ and $\phi=\frac{\pi}{6}$ according to the optimal RA deflection angles in (18). It is observed that since RA array can focus its radiation power through aligning the main lobes of all antenna elements towards the desired direction, it significantly improves the array gain and narrows the beamwidth as compared to the fixed-antenna array. This result indicates that our proposed RA-enabled system has the ability to improve the communication performance by reconfiguring the array directional gain pattern according to different wireless environments and applications.

### v4

@@ II-A Antenna Boresight Rotation
where $\theta_{\mathrm{max}}\in[0,\frac{\pi}{2}]$ is the maximum zenith angle that each RA is allowed to adjust. [^1] To characterize the fundamental performance of the proposed RA system, we assume that the orientation/boresight of each RA can be continuously tuned in its rotational range, subject to the constraint in (4), while in practice it is usually chosen from a finite number of available directions for the ease of hardware implementation. The design of RA systems with discrete orientation/boresight rotation will be left for future work.
@@ II-B Channel Model
The effective antenna gain for each RA depends on both the signal arrival/departure angle and antenna directional gain pattern. In this paper, we consider the following widely used directional gain pattern for each RA (corresponding to the directional antenna with one narrow main lobe and negligible side lobes) [27]
@@
where $(\epsilon,\varphi)$ is a pair of incident angles of the signal with respect to the RA’s current boresight direction as shown in Fig. 2(b), $p\geq 0$ is the directivity factor that characterizes the beamwidth of the antenna’s main lobe, and $G_{0}=2(2p+1)$ is the maximum gain in the boresight direction (i.e., $\epsilon=0$) that meets the law of power conservation.
We consider the scattering environment with $Q$ distributed scatterer clusters, where the position of scatterer cluster $q$ is represented by $\mathbf{c}_{q}\in\mathbb{R}^{3\times 1}$ with $q=1,2,\dots,Q$. In addition, we consider the narrow-band frequency-flat channel model for ease of exposition. [^2] The current narrow-band model can be extended to wideband scenarios [28, 29] by employing orthogonal frequency division multiplexing (OFDM) technology, which divides the wideband channel into multiple orthogonal sub-bands, each undergoing frequency-flat fading. For this extension, the beamforming and orientations/boresights of RAs need to be jointly optimized across all sub-bands to maximize their achievable sum rate. Based on the Friis Transmission Equation and the directional gain pattern adopted in (5), the channel power gain between user $k$ and RA $n$ can be modeled as [30]
$g_{\mathrm{u},k}(\vec{\mathbf{f}}_{n})$ $\approx\int_{\mathbb{A}_{n}}\frac{1}{4\pi\|\mathbf{u}_{k}-\mathbf{a}\|^{2}}G_{0}\left(\frac{\vec{\mathbf{f}}_{n}^{T}(\mathbf{u}_{k}-\mathbf{a})}{\|\mathbf{u}_{k}-\mathbf{a}\|}\right)^{2p}\mathrm{d}\mathbf{a}$
$=\frac{A}{4\pi r_{k,n}^{2}}G_{0}\operatorname{cos}^{2p}(\epsilon_{k,n}),$ (6)
@@
where the integral space $\mathbb{A}_{n}=\left[n_{x}\Delta-\frac{\sqrt{A}}{2},n_{x}\Delta+\frac{\sqrt{A}}{2}\right]\times\left[n_{y}\Delta-\frac{\sqrt{A}}{2},n_{y}\Delta+\frac{\sqrt{A}}{2}\right]$ corresponds to the surface region of RA $n$, $\mathbf{a}\in\mathbb{A}_{n}$ represents any point on plane $\mathbb{A}_{n}$, and $\operatorname{cos}(\epsilon_{k,n})\triangleq\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{u}}_{k,n}$ is the projection between $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{u}}_{k,n}$ with $\vec{\mathbf{u}}_{k,n}\triangleq\frac{\mathbf{u}_{k}-\mathbf{w}_{n}}{\|\mathbf{u}_{k}-\mathbf{w}_{n}\|}$ being the direction vector from RA $n$ to user $k$. Similarly, the channel power gain between scatterer cluster $q$ and RA $n$ is modeled as
$g_{\mathrm{c},q}(\vec{\mathbf{f}}_{n})\approx\frac{A}{4\pi d_{q,n}^{2}}G_{0}\operatorname{cos}^{2p}(\tilde{\epsilon}_{q,n}),$ (7)
@@
where $d_{q,n}=\|\mathbf{c}_{q}-\mathbf{w}_{n}\|$ is the distance between scatterer cluster $q$ and RA $n$, and $\operatorname{cos}(\tilde{\epsilon}_{q,n})\triangleq\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{c}}_{q,n}$ is the projection between $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{c}}_{q,n}$ with $\vec{\mathbf{c}}_{q,n}\triangleq\frac{\mathbf{c}_{q}-\mathbf{w}_{n}}{\|\mathbf{c}_{q}-\mathbf{w}_{n}\|}$ denoting the direction vector from RA $n$ to scatterer cluster $q$. Note that the channel power gains modeled in (6) and (7) account for both path gain and directional gain.
@@
For the multipath channel between RA $n$ and user $k$, by considering the geometric near-field propagation, the line-of-sight (LoS) channel component $h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{n})$ and the non-LoS (NLoS) channel component $h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{n})$ can be separately modeled by [31, 32]
$h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{n})$ $=\sqrt{g_{\mathrm{u},k}(\vec{\mathbf{f}}_{n})}e^{-j\frac{2\pi}{\lambda}r_{k,n}},$ (8)
$h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{n})$ $=\sum_{q=1}^{Q}{\frac{\sqrt{\sigma_{q}g_{\mathrm{c},q}(\vec{\mathbf{f}}_{n})}}{t_{k,q}}}e^{-j\frac{2\pi}{\lambda}(d_{q,n}+t_{k,q})+j\chi_{q}},$ (9)
@@
where $\lambda$ is the signal wavelength, $\sigma_{q}$ represents the radar cross section (RCS) of scatterer cluster $q$, $\chi_{q}$ represents the phase shift introduced by scatterer cluster $q$, and $t_{k,q}=\|\mathbf{u}_{k}-\mathbf{c}_{q}\|$ denotes the distance between user $k$ and scatterer cluster $q$. Thus, by superimposing the LoS and NLoS channel components, the overall multipath channel between user $k$ and the BS is given by
$\mathbf{h}_{k}(\mathbf{F})=\mathbf{h}_{k}^{\mathrm{LoS}}(\mathbf{F})+\mathbf{h}_{k}^{\mathrm{NLoS}}(\mathbf{F}),$ (10)
where $\mathbf{F}\triangleq[\vec{\mathbf{f}}_{1},\vec{\mathbf{f}}_{2},\dots,\vec{\mathbf{f}}_{N}]\in\mathbb{R}^{3\times N}$ is the pointing matrix of all RAs, $\mathbf{h}_{k}^{\mathrm{LoS}}(\mathbf{F})\triangleq[h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{1}),h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{2}),\dots,h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{N})]^{T}$ and $\mathbf{h}_{k}^{\mathrm{NLoS}}(\mathbf{F})\triangleq[h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{1}),h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{2}),\dots,h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{N})]^{T}$ are the LoS and NLoS channel components from the BS to user $k$, respectively. It is observed that variations in $\mathbf{F}$ lead to the rotation of the directional gain pattern of each RA, thereby altering the effective multipath channel in (10). Note that the current geometric channel model can also be extended to incorporate more realistic propagation characteristics by adopting the standardized path loss models defined in 3GPP TR 38.901 [33].
$\mathbf{y}=\sum_{k=1}^{K}{\mathbf{h}_{k}(\mathbf{F})\sqrt{P_{k}}s_{k}}+\mathbf{n},$ (11)
where $P_{k}$ and $s_{k}$ are the transmit power and information-bearing signal of user $k$, respectively, and $\mathbf{n}$ is the additive white Gaussian noise (AWGN) vector, following the zero-mean CSCG distribution with variance $\sigma^{2}$, i.e., $\mathbf{n}\sim\mathcal{N}_{c}(\mathbf{0},\sigma^{2}\mathbf{I}_{N})$. Upon receiving $\mathbf{y}$, the BS applies a linear receive beamforming vector $\mathbf{v}_{k}^{H}\in\mathbb{C}^{1\times N}$ with $\|\mathbf{v}_{k}\|=1$ to extract the signal of user $k$, i.e.,
$y_{k}=\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})\sqrt{P_{k}}s_{k}+\sum_{j\neq k}{\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})\sqrt{P_{j}}s_{j}}+\mathbf{v}_{k}^{H}\mathbf{n}.$ (12)
Accordingly, the receive SINR at the BS for decoding the information from user $k$ is given by
$\gamma_{k}=\frac{\bar{P}_{k}|\mathbf{v}_{k}^{H}\mathbf{h}_{k}(\mathbf{F})|^{2}}{\sum_{j\neq k}{\bar{P}_{j}|\mathbf{v}_{k}^{H}\mathbf{h}_{j}(\mathbf{F})|^{2}}+1},$ (13)
Remark 1: Although antenna orientation/boresight adjustments inevitably alter antenna polarization characteristics in practice, the proposed RA architecture introduces only minor variations in polarization characteristics due to its constrained rotational range at fixed antenna position and the absence of self-rotation. Moreover, the polarization ellipticity and handedness remain constant throughout the adjustment process, and thus efficient polarforming techniques [34, 35] can be applied to compensate the polarization induced channel variation such that the orientation/boresight adjustment by RA is still effective. Therefore, to focus on the effects of the directional gain pattern and explore the spatial DoFs offered by antenna orientation/boresight adjustments, polarization modeling is not considered in this paper for simplicity. Nevertheless, polarization modeling can be incorporated into future work on RA by integrating polarization matching efficiency into the channel power gain, following approaches similar to [34] and [35]. Such extensions would require joint consideration of both the directional gain pattern and polarization matching during orientation/boresight optimization.
@@ II-C Min-SINR Maximization Problem
In this paper, we aim to maximize the minimum SINR among all the users by jointly optimizing the receive beamforming matrix $\mathbf{V}\triangleq[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{K}]$ and RA’s pointing matrix $\mathbf{F}$, subject to the rotational constraint given in (4). Thus, the min-SINR maximization problem is formulated as
@@
s.t. $0\leq\operatorname{arccos}(\vec{\mathbf{f}}_{n}^{T}\mathbf{e}_{3})\leq\theta_{\mathrm{max}},\;\forall n,$ (14b)
$\|\vec{\mathbf{f}}_{n}\|=1,\;\forall n,$ (14c)
$\|\mathbf{v}_{k}\|=1,\;\forall k,$ (14d)
where constraint (14b) is equivalent to (4) for ensuring that the rotation of RAs’ boresight directions does not exceed the given range, and constraint (14c) ensures that $\vec{\mathbf{f}}_{n}$ is a unit vector.
In practice, channel estimation is necessary at the BS to obtain the required channel state information (CSI). Since adjusting the antenna’s orientation/boresight rather than its position does not change the propagation geometry, as shown in (8) and (9), conventional MIMO channel estimation methods for fixed antennas remain applicable to the RA system [22]. Furthermore, the flexible adjustments of RAs’ orientations/boresights enable the BS to capture pilot signals from diverse directions of arrival, yielding richer channel information and enhanced CSI estimation accuracy.
Denoting $\hat{\mathbf{h}}_{k}(\mathbf{F})$ as the estimated CSI of the link from user $k$ to the BS, the received signal is expressed as
$\mathbf{y}=\sum_{k=1}^{K}{\hat{\mathbf{h}}_{k}(\mathbf{F})\sqrt{P_{k}}s_{k}}+\sum_{k=1}^{K}{\check{\mathbf{h}}_{k}\sqrt{P_{k}}s_{k}}+\mathbf{n},$ (15)
where $\check{\mathbf{h}}_{k}\triangleq\mathbf{h}_{k}(\mathbf{F})-\hat{\mathbf{h}}_{k}(\mathbf{F})$ stands for the CSI estimation error that is uncorrelated to $\mathbf{h}_{k}(\mathbf{F})$. If the CSI estimation error follows the zero-mean complex Gaussian distribution, i.e., $\sum_{k=1}^{K}{\check{\mathbf{h}}_{k}\sqrt{P_{k}}s_{k}}\sim\mathcal{N}_{c}\left(\mathbf{0},e^{2}\mathbf{I}_{N}\right)$ with $e$ accounting for the level of channel estimation error [36], the resultant error term can be incorporated into the AWGN term in (15) for simplicity. Accordingly, the receive SINR for user $k$ is given by
$\bar{\gamma}_{k}=\frac{\tilde{P}_{k}|\mathbf{v}_{k}^{H}\hat{\mathbf{h}}_{k}(\mathbf{F})|^{2}}{\sum_{j\neq k}{\tilde{P}_{j}|\mathbf{v}_{k}^{H}\hat{\mathbf{h}}_{j}(\mathbf{F})|^{2}}+1},$ (16)
where $\tilde{P}_{k}\triangleq\frac{P_{k}}{\sigma^{2}+e^{2}}$ denotes user $k$’s effective transmit SNR. It should be pointed out that $\bar{\gamma}_{k}$ in (16) serves as a performance lower bound of the SINR modeled in (13) due to the estimation error. Additionally, by carefully examining the SINR in (16), we find that it has a form similar to the SINR given in (13). Therefore, the subsequent theoretical analysis and optimization algorithms based on problem (P1) remain applicable in practical scenarios considering CSI estimation errors.
@@ III Single-User Case with Free-Space Propagation
In this section, we consider the single-user and free-space propagation setup, i.e., $K=1$ and $Q=0$, to draw essential insights into (P1). Thus, the channel modeled in (10) reduces to $\mathbf{h}_{1}(\mathbf{F})=\mathbf{h}_{1}^{\mathrm{LoS}}(\mathbf{F})$. In this case, since no inter-user interference is present, problem (P1) is simplified to (by dropping the user index)
@@ III-A Optimal Closed-Form Solution
For any given RA’s pointing matrix $\mathbf{F}$ in the single-user case, it is known that the MRC beamformer is the optimal receive beamforming solution to problem (P2) [37], i.e., $\mathbf{v}_{\mathrm{MRC}}=\frac{\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})}{\|\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})\|}$. Thus, substituting $\mathbf{v}_{\mathrm{MRC}}$ into (17a) yields the following SNR expression,
$\gamma=\bar{P}\|\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})\|^{2}=\frac{\bar{P}A}{4\pi}\sum_{n=1}^{N}{\frac{G_{0}\operatorname{cos}^{2p}(\epsilon_{n})}{r_{n}^{2}}}.$ (18)
Exploiting the structure in (18), problem (P2) can be decomposed into $N$ subproblems, each of which independently optimizes the pointing vector of one RA. For RA $n$, the corresponding subproblem is given by
@@
s.t. $0\leq\operatorname{arccos}(\vec{\mathbf{f}}_{n}^{T}\mathbf{e}_{3})\leq\theta_{\mathrm{max}},$ (19b)
$\|\vec{\mathbf{f}}_{n}\|=1,$ (19c)
where the constant term is omitted in (19a). By maximizing the projection between the unit vector $\vec{\mathbf{f}}_{n}$ and $\vec{\mathbf{u}}_{n}$, the optimal solution to problem (P3) is obtained as
$\vec{\mathbf{f}}_{n}^{\star}=\left[\operatorname{sin}\theta_{\mathrm{z},n}^{\star}\operatorname{cos}\theta_{\mathrm{a},n}^{\star},\operatorname{sin}\theta_{\mathrm{z},n}^{\star}\operatorname{sin}\theta_{\mathrm{a},n}^{\star},\operatorname{cos}\theta_{\mathrm{z},n}^{\star}\right]^{T},$ (20)
where
$\theta_{\mathrm{z},n}^{\star}$ $=\mathrm{min}\left\{\operatorname{arccos}\left(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3}\right),\theta_{\mathrm{max}}\right\},$ (21a)
@@
with $\mathbf{e}_{1}\triangleq[1,0,0]^{T}$, $\mathbf{e}_{2}\triangleq[0,1,0]^{T}$, and $\mathbf{e}_{3}\triangleq[0,0,1]^{T}$.
According to the optimal pointing vector in (20), it can be inferred that each RA prefers to tune its antenna orientation/boresight towards the user. This is expected since the BS can achieve the maximum directional gain $NG_{0}$ when the boresight direction of each RA is aligned with the user direction, i.e., $\vec{\mathbf{f}}_{n}=\vec{\mathbf{u}}_{n}$.